OpenStudy (anonymous):

Which expression is equivalent to (4 + 7i)(3 + 4i)?

OpenStudy (anonymous):

28i^2+37i+12

OpenStudy (anonymous):

Which expression is equivalent to (4 + 7i)(3 + 4i)? –16 + 37i 12 – 28i 16 – 37i 37 + 16i

OpenStudy (anonymous):

i^2 is -1 so 28i^2 is just -28, so now your equation is 37i-16

OpenStudy (anonymous):

Use FOIL or whatever method you use to multiply two binomials. You get, 12 + 16i + 21i + 28i^2 = 12 + 37i + 28i^2 BUT, i^2 = -1, so you get 12 + 37i + 28(-1) = 12 + 37i - 28 = -16 + 37i

OpenStudy (anonymous):

Yeah me and easyaspi got the same answer so it's probably right.

OpenStudy (anonymous):

can anyone help me?

OpenStudy (anonymous):

Choice (1) is correct. Most important is to see a step-by-step solution so you can simplify it yourself if there were to be no choices.

OpenStudy (anonymous):

Which values of x and y would make the following expression represent a real number? (4 + 5i)(x + yi) x = 4, y = 5 x = –4, y = 0 x = 4, y = –5 x = 0, y = 5